1.     The Kilmarnock Campus tug of war team pulls with a force of 400N while the Ayr Campus team pulls with a 300N force.  What is the size of the resultant force and which team wins?


1.     A car (mass 1000kg) engine produces a force of 2000N.  There is a frictional force of 100N acting against the car (air resistance is negligible).  Calculate the resultant force and hence the cars acceleration.



2.     A parachutist and his parachute have a total mass of 100kg. At a certain point air resistance is measured as 400N

a.     Calculate the combined weight?

b.     Determine the resultant force & direction.

c.      Calculate the acceleration.



1
Expert's answer
2021-04-13T06:32:01-0400

Answer 1

Resultant force

F=400-300=100N

So Kilmarnock Campus tug of war team will win.

Answer 2

Resultant force

R=2000-100=1900N

Acceleration

"a=\frac{1900}{1000}=1.9m/s^2"


Answer 3

a)

Combine weight

"W=(100\times9.81)=981N"


b) resultant force

"F=981-400=581N\n(Downward)"

c)

Acceleration

"a=\frac{581}{100}=5.81m/s^2"